OpenStudy (anonymous):

Use the slope intercept form to graph the equation. 5x+y=0

OpenStudy (anonymous):

if you rearrange the formula you can get it into y=mx+b form which will be y=-5x+o which means your line will intersect at (0,0) and the slope is -5 which means it will go 5 units left and 1 unit up (-5/1).

OpenStudy (anonymous):

The slope intercept form is y=mx+b. So rewrite 5x+y=0.... y=-5x + 0. Now, plot a point at the y-intercept, or b, or 0 in this case. Then move down 5, and to the right one. (the slope is m, and in this case it is -5/1). Plot a point each time you move down 5, and to the right one. Continue this from each point you plot.

OpenStudy (anonymous):

omg yes sorry for saying the wrong this for the slope its rise over run so down 5 1 right not the way i said it